# SERN: Simulation-Enhanced Realistic Navigation for Multi-Agent Robotic Systems in Contested Environments

### v1

@@ III-E Environmental Perception and Navigation Planning
For navigation, 3D lidar is first transformed into a 2D scan by applying a condition: if an object within the lidar’s detection range surpasses a certain threshold (i.e., considered to be in close proximity to the robot), it is flagged as an obstacle. This transformation reduces computational complexity and enables the robot to perform path planning using the 2D representation.
@@
Here, $L_{3D}(r,\theta,z)$ represents the 3D lidar points, and $S_{2D}(r,\theta)$ represents the 3D lidar points. Simultaneous Localization and Mapping (SLAM) [36] is implemented on the 2D scan for navigation, where the robot computes a map and identifies an optimal path toward a final goal position.
To handle semantic segmentation, we developed a domain adaptation-based approach [37], leveraging 3D lidar data to enhance generalization across different domains (e.g., different environments). Domain adaptation is crucial for improving the model’s ability to generalize without requiring vast amounts of labeled data from the target domain. By aligning the source and target domain distributions, our model is able to better understand and segment the 3D lidar data in real-time, ensuring accurate perception (Fig. 5).
@@ IV Experiments and Results
This section details the implementation of our SERN framework, experimental scenario design, compares it with traditional ROS architecture on computational and communication metrics, and demonstrates its application in LiDAR point cloud-based semantic segmentation and navigation.
@@ IV-A Experimental Setup
To assess the effectiveness of our distributed SERN framework, we designed a test environment that simulates geographically separated ROS networks connected through the AuroraXR server. Fig. 2 illustrates the network topology used in our experiments, comprising two distinct ROS networks, each operating on separate subnets. Each network includes physical robots, virtual robots, a ROS Master node, a ROS Bridge interfacing with the AuroraXR server, a WiFi access point for local wireless communication, and a network switch for wired connections.
Figure 6: Performance Comparison of AuroraXR ROS Bridge integrated SERN framework vs. Traditional ROS. We implemented the identical process of increasing the robot numbers by 2, 3 and 5 and recorded the CPU, memory usage, power consumption, and latency overhead of the AuroraXR server and traditional ROS Master. We have chosen the point cloud data generated by each of the robots as data modality to be read and visualized and illustrated the averaged values after 20 iterations of each setup.
The AuroraXR server acts as the central component bridging these two networks, facilitating communication between the geographically separated ROS environments. It connects to both networks via distinct network interfaces, enabling efficient routing of messages between the ROS nodes in each network. This setup allows us to rigorously test various scenarios that are crucial for distributed multi-agent systems, including real-time synchronization of robot movements across networks, transmission of sensor data between robots, and distributed control where operators on one network can command robots on the other. Additionally, we simulated latency and network disruptions to evaluate the system’s robustness under adverse conditions.
We have selected an experimental scenario to solidify our conceptualization of SERN, which involves real-time navigation, coordination with the virtual environment, and the application of an on-the-fly semantic segmentation technique. We have created an identical virtual representation of an area of the UMBC CARDS Lab [38] using Unity, where we have placed a car as an obstacle (Fig. 5). The robot’s task is to avoid it and navigate through a wooden bridge, provided it is not obstructed by other adversaries. To better understand the virtual-physical coordination, we have maintained some discrepancies between the two environments. For example, the virtual scenario includes fallen tree logs on the bridge, while the actual environment’s bridge remains navigable. To initiate the experiment, the physical robot will continuously gather environmental information from the virtual robot, all while operating simultaneously. However, the physical robot should also detect any discrepancies among the virtual information and act using its own semantic segmentation algorithm. Information exchange is primarily carried out via ROStopics and transmitted over the AuroraXR server. Additionally, an app proxy is set up within each agent. Extending the experimental scenario, we systematically augmented the quantity of topics originating from several robots and Rosbags, and subsequently assessed the computational and communication metrics.
@@ IV-B Performance Analysis
The synchronization between physical and virtual robots was highly accurate, with an average positional error of less than 5cm and rotational error under 2 degrees. This high fidelity in synchronization ensures that virtual simulations closely represent real-world scenarios, enhancing the reliability of mission planning and training exercises. The cross-reality integration showed promising results, with virtual object insertions reflected in the physical robot’s perception with an average positional error of only 2.3cm. The system achieved a 10Hz refresh rate for physical-to-virtual environment updates, allowing for near real-time synchronization. This capability enhances situational awareness and enables more effective remote operation and mission planning.
To evaluate the scalability and performance of the AuroraXR integration, we conducted a series of tests comparing our system against a traditional ROS setup across different numbers of agents. The metrics considered include CPU usage, memory usage, power consumption, and the latency for sending commands and data as the number of agents increases.
@@
As shown in Fig. 6, the AuroraXR ROS Bridge demonstrates superior scalability compared to the traditional ROS setup across an increasing number of agents. The AuroraXR Bridge consistently exhibits lower CPU and memory usage, as well as reduced power consumption, which are critical factors for deploying distributed robotic systems in resource-constrained environments. Notably, the power consumption overhead on the server node shows an improvement of a range from approximately 10% to 39% while increasing the number of robotic agents from 2 to 5, compared to traditional ROS Master, mainly due to the offloading technique adopted in the SERN architecture. Furthermore, the latency for both command and data transmission remains significantly lower with AuroraXR, with improvements ranging from 15.7% to 24.7% as the number of agents increases. This indicates that the AuroraXR integration not only enhances communication efficiency but also scales effectively, maintaining performance advantages even as the system load grows.
The reduced CPU consumption in SERN is attributed to its intelligent communication management and decoupled architecture, which effectively minimizes the processing burden on individual nodes. This is particularly advantageous when the quantity of robots grows, as conventional ROS has challenges with increased CPU requirements owing to its synchronous and more tightly coupled communication structures. The SERN design facilitates the transfer and enhancement of data processing amongst several components, therefore enabling more effective resource management, subsequently improving memory utilization and power consumption. The utilization of ZMQ for serialized binary data transfer improves the latency values while handling data originating from several robots by reducing the overhead caused by the consistently growing ROStopics commonly observed in conventional ROS communication stacks.
@@
Dynamic Configuration Validation: Our real robot testing setup demonstrates the system’s dynamic configuration capability. Initially, multiple robots are connected to the bridge on one network, with visualizers on another. The ROS bridge, set to dynamic topic discovery mode, continuously monitors for new nodes and topics. When a new robot joins the network, the bridge automatically detects it and subscribes to its relevant topics, integrating its data into the existing flow without system interruption. This approach enables seamless expansion of the robot team during operation, enhancing flexibility and scalability.
@@
Cross-Network ROS Communication Testing: To validate the effectiveness of the SERN framework in managing cross-network communication, we conducted tests using ROS publisher and subscriber nodes connected via the AuroraXR server. In this setup, each ROS bridge was deployed on separate laptops operating on distinct networks, simulating a distributed multi-agent environment. On the publishing side, ROS traffic was generated using ROS bag files, which emulates real-world sensor data and control commands. This data was published to the ROS bridge, which subscribes to the incoming messages. The bridge then encapsulates the data within a custom envelope format optimized for transmission across the AuroraXR server. On the receiving side, the ROS bridge re-publishes the data to the ROS broadcast network, where it is subscribed to by various nodes, including a data visualizer for monitoring. This end-to-end test setup demonstrates the bi-directional communication capabilities of SERN, allowing for seamless data flow and interaction between distributed ROS networks. A video demonstration is provided. [^1] Video demonstration available at: https://youtu.be/BqNoqDPqhNc
@@ V Conclusion, Limitations, and Future Directions
In this paper, we introduced SERN, a framework that bridges physical and virtual environments to enhance coordination and decision-making in multi-agent robotic systems. Testing in simulated and real-world scenarios showed improved synchronization and navigation performance. However, SERN’s reliance on network stability can impact performance in disrupted or low-connectivity environments, and its real-time processing poses challenges when processing multi-modal data for a common mission with increasing agent numbers. Future work will focus on enhancing resilience to network disruptions, exploring edge computing to reduce latency, enhancing the capability of SERN for remote robotics experimentation across different geolocations, and expanding support for heterogeneous robot teams, including aerial and underwater vehicles, with advanced machine learning for predictive navigation.

### v2

@@ III-F Environmental Perception and Navigation Planning
To evaluate command and control, we also integrated semantic segmentation and navigation capabilities in both physical and virtual domains, enabling simultaneous command execution. For navigation, 3D LiDAR is first transformed into a 2D scan by applying a condition: if an object within the LiDAR detection range surpasses a certain threshold (i.e., considered to be in close proximity to the robot), it is flagged as an obstacle. This transformation reduces computational complexity and enables the robot to perform path planning using the 2D representation.
@@
Here, $L_{3D}(r,\theta,z)$ represents the 3D lidar points, and $S_{2D}(r,\theta)$ represents the 3D lidar points. Simultaneous Localization and Mapping (SLAM) is implemented on the 2D scan for navigation, where the robot computes a map and identifies an optimal path toward a final goal position. To handle semantic segmentation, we developed a domain adaptation-based approach [40], leveraging 3D lidar data to enhance generalization across different domains (e.g., different environments). Domain adaptation is crucial for improving the model’s ability to generalize without requiring vast amounts of labeled data from the target domain. By aligning the source and target domain distributions, our model is able to better understand and segment the 3D lidar data in real-time, ensuring accurate perception (Fig. 5).
@@ IV Experiments and Results
This section details the implementation of our SERN framework, experimental scenario design, compares it with traditional ROS architecture.
@@ IV-A Experimental Setup
To assess the effectiveness of our distributed SERN framework, we designed a test environment that simulates geographically separated ROS networks connected through the AuroraXR server. Fig. 4 illustrates the network topology used in our experiments, comprising two distinct ROS networks, each operating on separate subnets. Each network includes physical robots, virtual robots, a ROS Master node, a ROS Bridge interfacing with the AuroraXR server, a WiFi access point for local wireless communication, and a network switch for wired connections. The AuroraXR server bridges the networks through distinct interfaces, routing messages between ROS nodes across geographical boundaries. This architecture enables testing of critical distributed capabilities: cross-network robot synchronization, sensor data exchange, and remote operator control across network domains.
Figure 6: Performance Comparison of SERN ROS Bridge framework vs. Traditional ROS. We implemented the identical process of increasing the robot numbers by 2, 3 and 5 and recorded the CPU, memory usage, power consumption, and latency overhead of the AuroraXR server (integrated with SERN) and traditional ROS Master. We have chosen the point cloud data generated by each of the robots as data modality to be read and visualized and illustrated the averaged values after 20 iterations of each setup.
We have selected an experimental scenario to solidify our conceptualization of SERN, which involves real-time navigation, coordination with the virtual environment, and the application of an on-the-fly semantic segmentation technique. We have created an identical virtual representation of an area of the UMBC CARDS Lab [41] using Unity, where we have placed a car as an obstacle (Fig. 5). The robot’s task is to avoid it and navigate through a wooden bridge, provided it is not obstructed by other adversaries. To better understand the virtual-physical coordination, we have maintained some discrepancies between the two environments. For example, the virtual scenario includes fallen tree logs on the bridge, while the actual environment’s bridge remains navigable. To initiate the experiment, the physical robot will continuously gather environmental information from the virtual robot, all while operating simultaneously. However, the physical robot should also detect any discrepancies among the virtual information and act using its own semantic segmentation algorithm. Information exchange is primarily carried out via ROStopics and transmitted over the AuroraXR server. Extending the experimental scenario, we systematically augmented the quantity of topics originating from several robots and Rosbags, and subsequently assessed the computational and communication metrics.
@@ IV-B Performance Analysis
The synchronization between physical and virtual robots was highly accurate, with an average positional error of less than 5cm and rotational error under $2^{\circ}$. This high fidelity in synchronization ensures that virtual simulations closely represent real-world scenarios, enhancing the reliability of mission planning and training exercises. The cross-reality integration showed promising results, with virtual object insertions reflected in the physical robot’s perception with an average positional error of only 2.3cm. The system achieved a 10Hz refresh rate for physical-to-virtual environment updates, allowing for near real-time synchronization. This capability enhances situational awareness and enables more effective remote operation and mission planning.
@@
To evaluate the scalability and performance of the SERN integration, we conducted a series of tests comparing our system against a traditional ROS setup across different numbers of agents. The metrics considered include CPU usage, memory usage, power consumption, and the latency for sending commands and data as the number of agents increases. As shown in Fig. 6, the our method demonstrates superior scalability compared to the traditional ROS setup across an increasing number of agents. The SERN Bridge consistently exhibits lower CPU and memory usage, as well as reduced power consumption, which are critical factors for deploying distributed robotic systems in resource-constrained environments. Notably, the power consumption overhead on the server node shows an improvement of a range from approximately 10% to 39% while increasing the number of robotic agents from 2 to 5, compared to traditional ROS Master, mainly due to the offloading technique adopted in the SERN architecture. Furthermore, the latency for both command and data transmission remains significantly lower with SERN, with improvements ranging from 15.7% to 24.7% as the number of agents increases. This indicates that the SERN integration not only enhances communication efficiency but also scales effectively, maintaining performance advantages even as the system load grows.
@@
Dynamic Configuration Validation: Our real robot testing setup demonstrates the system’s dynamic configuration capability. Initially, multiple robots are connected to the bridge on one network, with visualizers on another. The SERN ROS bridge, set to dynamic topic discovery mode, continuously monitors for new nodes and topics. When a new robot joins the network, the bridge automatically detects it and subscribes to its relevant topics, integrating its data into the existing flow without system interruption. This approach enables seamless expansion of the robot team during operation, enhancing flexibility and scalability.
@@
Cross-Network ROS Communication Testing: To validate the effectiveness of the SERN framework in managing cross-network communication, we conducted tests using ROS publisher and subscriber nodes connected via the AuroraXR server. In this setup, each SERN ROS bridge was deployed on separate laptops operating on distinct networks, simulating a distributed multi-agent environment. On the publishing side, ROS traffic was generated using ROS bag files, which emulates real-world sensor data and control commands. This data was published to the SERN ROS bridge, which subscribes to the incoming messages. The bridge then encapsulates the data within a custom envelope format optimized for transmission across the AuroraXR server. On the receiving side, the SERN ROS bridge re-publishes the data to the ROS broadcast network, where it is subscribed to by various nodes, including a data visualizer for monitoring. This end-to-end test setup demonstrates the bi-directional communication capabilities of SERN, allowing for seamless data flow and interaction between distributed ROS networks. A video demonstration is provided in our project page. [^1] Video demonstration available at: https://pralgomathic.github.io/sern.multi-agent/
@@ V Conclusion, Limitations, and Future Directions
In this paper, we introduced SERN, a framework that bridges physical and virtual environments to enhance coordination and decision-making in multi-agent robotic systems. Testing in simulated and real-world scenarios showed improved synchronization and navigation performance. However, SERN’s reliance on network stability can impact performance in disrupted or low-connectivity environments, and its real-time processing poses challenges when processing multi-modal data for a common mission with increasing agent numbers. Although our experiments included a combination of physical and virtual agents—a single robot and numerous virtual agents—the testing of scalability in actual environments is still lacking. Better validation and multi-agent cooperation in varied contexts will be achieved in subsequent research by adding more physical robots.